I'm sturggling with this problem. Any help would be greatly appreciated.
Solve the following non-linear system by substitution.
x^2+y^2=16
x-2y=0

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(57224)

Solve the following non-linear system by substitution.
x^2+y^2=16
x-2y=0
-----------
x = 2y
Substitute into the 1st equation to solve for "y":
(2y)^2 + y^2 = 16
4y^2 + y^2 = 16
5y^2 = 16
y^2 = 16/5
------------
y = 4/sqrt(5) = (4/5)sqrt(5)
OR
y = (-4/5)sqrt5
-----------------
Substitute into x = 2y to get:
x = 2[(4/5)sqrt5] = (8/5)sqrt(5)
OR
x = 2(-4/5)sqrt5 = (-8/5)sqrt(5)
=====================
Cheers,
Stan H.

Answer by checkley71(8403)

x-2y=0 or x=2y
now substitute 2y for x in the other equation & solve for y.
(2y)^2+y^2=16
4y^2+y^2=16
5y^2=16
y^2=16/5
y=4/sqrt5 answer.
x-2(4/sqrt5)=0
x-8/sqrt5=0
x=8/sqrt5 answer.